OpenStudy (anonymous):

Simplify: x^2-20x+14 ----------- x+3 what is the integer remainder?

OpenStudy (anonymous):

use the remainder theorem to do this; we are dividing by x+3 so we put x = -3 into f(x) - ths gives the remainder (-3)^2 - 20(-3) + 14 = 9 +60 + 14 = 83 remainder is 83

OpenStudy (anonymous):

any more questions ?